# Fokker-Planck approach for thermal fluctuations in antiferromagnetic systems

### v1

@@ III Equations of motion
In this Section, starting from the FP equation shown in Eq. (10), we derive the equation of motion for the spin polarization, identified as the first moment garanin1997fokker, and for the two-time spin-spin correlation function likharev2019statistical; kogan2008electronic.
@@ III.1 Spin polarization
where on the right-hand side the first line represents the precession term, the second the diffusion term, and the third the Gilbert damping term. For the sake of clarity, we exclude the inter-sublattice Gilbert term $\Lambda^{\prime}$ from the present discussion. The full derivation of the equation of motion of the spin polarization, including $\Lambda^{\prime}$, is provided in Appendix A. Note that the anisotropy terms $\delta D_{\rm T}$ and $\delta\lambda$ only affect the in-plane component of the magnetization, and that the evaluation of the time derivative of the first-order moment involves higher-order moments (namely, second and third). The difficulty of solving the FP equation is transferred to the generation of an infinite hierarchical set of equations, which involves the higher-order moments. To solve this, in analogy to the approach used in garanin1997fokker for a ferromagnetic classical model, we apply the MF approximation. The PDF $\mathcal{P}_{\mathrm{MF}}$ assumes that the spins of the sublattices, ${\bf\it s}_{A}$ and ${\bf\it s}_{B}$, are independent. Each one is coupled to an effective external field $k_{\rm B}T{\bf\it\xi}_{\mathrm{A}}$ and $k_{\rm B}T{\bf\it\xi}_{\mathrm{B}}$, respectively, where $k_{\rm B}T$ is the thermal energy  garanin1997fokker. The PDF $\mathcal{P}_{\mathrm{MF}}$ is expressed as
@@
where $B(\xi)$ denotes the Langevin function langevin1905
@@
$\dot{\bm{m}}_{\alpha}$ $=\bm{m}_{\alpha}\times\expectationvalue{{\bm{h}}_{\alpha}}_{\rm MF}$
@@
Eq. (17) coincides with the nonlinear equation that determines the global minimum of the free energy via the classical Hamiltonian, as expressed in Eq. (1), within the MF approximation goldenfeld_book. Proceeding with the assumption that there is a stationary solution with spins lying in the $x$-$y$ plane, e.g. along the $x$ direction such that ${\bf\it\xi}_{\mathrm{A}}=-{\bf\it\xi}_{\mathrm{B}}=\xi{\bf\it u}_{x}$, the resulting expression for $\xi$ is given by
@@ III.2 Spin-spin correlation functions
where the indices $i$ and $\ell$ refer to the Cartesian coordinates $x,y,z$, $\mathcal{P}({\bf\it\mathcal{S}},t_{1}|{\bf\it\mathcal{S}}^{\prime},t_{0})$ is the conditional probability of having the spin configuration ${\bf\it\mathcal{S}}=({\bf\it s}_{\mathrm{A}},{\bf\it s}_{\mathrm{B}})$ at time $t_{1}$ assuming that the configuration at time $t_{0}$ was ${\bf\it\mathcal{S}}^{\prime}=({\bf\it s}^{\prime}_{\mathrm{A}},{\bf\it s}^{\prime}_{\mathrm{B}})$, $P_{0}({\bf\it\mathcal{S}}^{\prime})$ is the stationary probability distribution likharev2019statistical. Since we enforce the AFM condition $\bm{m}_{A}=-\bm{m}_{B}$ with $\Delta>0$, and we set ${\bf\it\xi}_{\mathrm{A}}=-{\bf\it\xi}_{\mathrm{B}}=\xi{\bf\it u}_{z}$. In the stationary regime, the correlation functions depend only on the time difference, so that $g_{\alpha i,\beta\ell}(t_{1},t_{0})\to g_{\alpha i,\beta\ell}(t)$, with $t=t_{1}-t_{0}$ for simplicity.
By differentiating Eq. (23) with respect to time through the use of Eq. (10), we derive the equation governing the dynamics of the correlation functions. The complete derivation is detailed in Appendix B, where it is shown that achieving an exact resolution of the motion equation requires solving a hierarchical series of equations. Since this hierarchy is infinite, an exact solution is unfeasible. Therefore, to derive a closed expression for the correlation functions, it becomes necessary to truncate the hierarchy. In the spirit of the Gaussian approximation, the truncation is performed by assuming that Wick’s theorem is applicable to the correlation functions of fluctuations of order higher than second goldenfeld_book, allowing them to be expressed in terms of second-order correlation functions. Initial conditions are fixed considering that in the limit $t\to 0$, we have $\mathcal{P}({\bf\it\mathcal{S}},t|{\bf\it\mathcal{S}}^{\prime},0)\rightarrow\delta({\bf\it\mathcal{S}}-{\bf\it\mathcal{S}}^{\prime})$. It follows that
@@ III.2.2 In-plane dynamics
which is expressed in terms of the spin-wave energy gap at zero-temperature $\Omega_{0}=\sqrt{\Delta(2J+\Delta)}$ rezende2019introduction. For the sake of clarity, the inter-sublattice terms $\lambda^{\prime}$ and $\delta\lambda^{\prime}$ have been omitted here. They have been examined in detail in Appendix B.2, where it is shown how they are incorporated into $\Gamma_{\bot}$ and $\Gamma_{\bot}^{\prime}$.
@@ IV Applications
Employing the method used to derive the equations of motion for spin polarization and spin-spin correlation functions, we examine two potential applications. Firstly, we investigate the dynamics of the spin-wave generated within the AFM phase kamra_prb_2018. Secondly, we show a phenomenological model tailored to describe the effects of magnetic fluctuations on the experimentally observed resistance fluctuations in MPX3 ghosh2024electronic.
@@ IV.1 Spin-waves
Starting from the results of Sec. III.1, we derive the equation of motion for spin polarization in the AFM configuration being perturbed by a weak, external magnetic field $\bm{h}^{\rm ext}$ to both sublattices gekht1976magnetic.
@@ IV.2 Thermal fluctuations in the resistance
As mentioned in the introduction, this work is motivated by the experimental observations presented in ghosh2024electronic, where vertical resistance fluctuations in a 2D antiferromagnetic semiconductor showed an anomalous behavior close below the Néel temperature. The compound used in the experiment is a transition metal-phosphorus-trichalcogenides (MPX3chittari2016electronic; nitschke2023valence; ghosh2024electronic; samal2021two. In these compounds, transition metal atoms (M) lay on a plane forming a honeycomb lattice, with negligible orbital overlap that inhibits conduction electrons from hopping within the plane, electron transport, however, takes place principally via the phosphorus-tricalchogenide groups (PX3) that take the shape of pyramids above and below the metal layer, and mediate longitudinal charge transport chittari2016electronic; nitschke2023valence; ghosh2024electronic; samal2021two and the sketch in Fig. 1.
@@
In this framework, one can approximate the electrons responsible for conduction as distinct from those responsible for the magnetic properties YamadaTakada1973; chittari2016electronic; nitschke2023valence. Based on the definition of the internal magnetic field for the sublattice in Eq. (3), and in the macrospin approximation, we can define the net internal magnetic field as the sum of the two fields ${\bf\it h}_{A}$ and ${\bf\it h}_{B}$:
@@
In the presence of the exact AFM configuration $\bm{s}_{A}=-\bm{s}_{B}$, the conducting electrons are not affected by a net internal magnetic field. When a disturbance disrupts the perfect AFM order, a non-zero net internal field can be produced. Using the methods shown in Sec. III, we develop a phenomenological model to describe thermal fluctuations as the source of this disturbance. In the limit of a weak external magnetic field, in conducting systems, electrical resistance typically depends quadratically on the intensity of the external field hu2008classical. According to the scheme represented in Fig. 1, conducting electrons experiencing a (stray) magnetic field, it is assumed that their transport generates an electrical resistance that depends quadratically on the intensity of the internal net magnetic field $\bm{h}$ as
$R=R_{0}+\frac{\kappa_{R}}{2}h^{2}~,$ (47)
where $R_{0}$ and $\kappa_{R}$ represent two phenomenological terms, $R_{0}$ is the resistance in the absence of interaction with the AFM system, and $\kappa_{R}$ is a proportionality factor assumed to be independent of the magnetic configuration. With the application of the FP equation, resistance $R(t)$ is considered a function dependent on random variables $\bm{s}_{\alpha}(t)$, linked by the internal net magnetic field $\bm{h}(t)$. It is useful to define the effective magnetoresistance as $R(t)=R_{0}+\kappa_{R}\quantity[R_{\parallel}(t)+R_{\bot}(t)]$, where
@@
$\expectationvalue{R}=R_{0}-\kappa_{R}\quantity[\Delta(2J+\Delta)\frac{B(\xi)}{\xi}+\frac{(J+\Delta)^{2}}{2}B^{2}(\xi)]~,$ (49)
where the terms in square brackets are the contributions in- and out-of-plane, respectively. Since we are interested in the resistance correlation function $\expectationvalue{\delta R(t)\,\delta R(0)}$, we write $\delta R(t)=R(t)-\expectationvalue{R}=\delta R_{\parallel}(t)+\delta R_{\bot}(t)$, where
@@
Our goal is to calculate the resistance power spectrum wiener1930generalized; khintchine1934korrelationstheorie; clerk2010introduction, defined as
@@
Fig. 2 (a) shows the out-of-plane component $S_{\parallel}(\omega)$ as a function of the temperature, covering the range from zero to Néel temperature, and setting $\Delta/J=0.35$. In particular, we consider three values of $\omega$: $\omega=\lambda J/10$ (solid red line), $\omega=\lambda J$ (solid green line) and $\omega=20\lambda J$ (solid blue line). Moreover, for each $\omega$, as reference, the dashed line and the dotted line represent, respectively, the contributions in Eq. (54) that involve the damping rate $\Gamma_{\parallel}^{\prime}$, which is $\propto 4\sigma_{z}^{2}B^{2}(\xi){\cal L}(\omega,\Gamma^{\prime}_{\parallel})+\sigma_{z}^{4}{\cal L}(\omega,2\Gamma_{\parallel}^{\prime})$ and $\Gamma_{\parallel}$, which is $\propto\sigma_{z}^{4}{\cal L}(\omega,2\Gamma_{\parallel})$. Here, we observe that in $S_{\parallel}(\omega)$, the term involving $\Gamma_{\parallel}$ (dotted lines) exhibits a monotonic behavior with temperature. This trend is intuite, as it arises directly from the increase in thermal fluctuations associated with a rise in temperature. Specifically, as the temperature increases, there is an enhancement of the prefactor $\sigma_{z}^{4}$. On the other hand, the terms of $S_{\parallel}(\omega)$ involving $\Gamma_{\parallel}^{\prime}$ exhibit a peak at a temperature lower than $T_{\rm N}$, and the specific value depends on the chosen $\omega$. In particular, unlike $\sigma_{z}$, the staggered magnetization decreases with increasing temperature, thus the interplay between $\sigma_{z}$ and $B(\xi)$ results in the non-monotonic behavior and the formation of a peak. Moreover, it has been observed that when $\omega\lesssim\lambda J$, the terms involving $\Gamma_{\parallel}^{\prime}$ play a dominant role. As a result, the entire $S_{\parallel}(\omega)$ displays a distinct peak at a temperature lower than $T_{\rm N}$. Fig. 2 (b) shows the behaviour of the characteristic damping rates $\Gamma_{\parallel}$ (black solid line) and $\Gamma_{\parallel}^{\prime}$ (black dashed line) as functions of temperature. The horizontal lines represent the values of $\omega$: $\lambda J/10$ and $\lambda J$, used in Fig. 2 (a), serving as references. Within the explored temperature range, $\Gamma_{\parallel}^{\prime}$ vanishes at Néel temperature, while $\Gamma_{\parallel}$ remains finite. This panel provides a quantitative criterion for determining when the term associated with $\Gamma_{\parallel}$ becomes negligible, that is, for $\omega<\Gamma_{\bot}$.
@@
Figure 3: a) In-plane term $S_{\bot}(\omega)$ of the resistance power spectrum, defined in Eq. (58), shown as a function of the temperature between the zero temperature and the Néel temperature, for different values of $\omega$: $\omega=\lambda J/10$ (red lines), $\omega=\lambda J$ (green lines), and $\omega=20\lambda J$ (blue lines). Here, the dimensionless parameter $\delta\lambda$ is set $\delta\lambda=\lambda$ in solid lines, $\delta\lambda=0$ in dashed lines, and $\delta\lambda=-\lambda/2$ in dash-dotted lines. Only a curve shows a non monotonic behavior, and it is the case with $\omega=20\lambda J$ and $\delta\lambda=\lambda$. b) $T^{2}\Gamma_{\bot}$ as a function of temperature, where the damping rate is defined in Eq. (30a). The association between line style and $\delta\lambda$ remains consistent with that in panel a), while the gray dotted line represents $\delta\lambda=\lambda/2$. Results in all panels have been obtained by setting $\Delta=0.35J$.
@@
which is proportional to the Lorentzian function centered at zero and of width $2\Gamma_{\bot}$ multiplied by $\sigma_{\bot}^{4}$. Since we are interested in temperatures below $T_{\rm N}$, we can replace $\sigma_{\bot}^{4}=(k_{\rm B}T)^{2}/(J+\Delta)^{2}$. Using Eq. (30a), we see that considering $\omega$ in units of $\lambda J$ makes $S_{\bot}(\omega)$ in the form of a product of $1/\lambda$ and a function of $\omega/(\lambda J)$ and $\delta\lambda/\lambda$. Again, this allows the results to be shown without explicit dependence on $\lambda$.
@@
Recalling that $\delta\lambda\neq 0$ implies that both the diffusion coefficients in Eq. (9b) and the Gilbert damping parameters are not identical, we note that for $\omega\lesssim\lambda J$ when $\delta\lambda>0$, the diffusion term and damping in planar spin-dynamics become more efficient compared to the cases with $1\leq\delta\lambda/\lambda<0$. The lower bound of $\delta\lambda/\lambda=-1$ ensures that diffusive coefficients and the Gilbert damping terms remain non-negative. Except for the case with $\omega=20\lambda J$ and $\delta\lambda=-1$, which will be discussed below, the in-plane contribution to the resistance power spectrum exhibits a monotonic temperature dependence, as expected. Here, increasing temperature improves fluctuations, which in turn leads to an increase in the prefactor $\sigma_{\bot}^{4}\propto T^{2}$.
In Fig. 3 (a), this case is shown with the solid blue line that displays a peak at a temperature slightly below the Néel temperature, located with a black triangle. The reason for this behaviour lies in the relation between $\omega$ and $\Gamma_{\bot}$. In particular, if $\omega\gg\Gamma_{\bot}$ across the entire temperature range examined, it is reasonable to approximate ${\cal L}(\omega,2\Gamma_{\bot})\approx\Gamma_{\bot}/(\pi\omega^{2})$, as a consequence of Eq. (58) one has $S_{\bot}(\omega)\propto T^{2}\Gamma_{\bot}/\omega^{2}$. The quantity in the numerator is shown in Fig. 3 b), where the different line styles are encoded in the same way as in Fig. 3 a), where we see that for $\delta\lambda/\lambda=1$ (solid line) $T^{2}\Gamma_{\bot}$ exhibits a peak below $T_{\rm N}$. Using Eqs. (21) and (30a), one finds that as $T\to 0$, the quantity $T^{2}\Gamma_{\bot}$ exhibits an increasing quadratic temperature dependence, since $\Gamma_{\bot}\approx(\lambda+\delta\lambda)(J+\Delta)$. On the other hand, as $T\to T_{\rm N}^{-}$, by using $\xi\propto\sqrt{1-T/T_{\rm N}}$, it shows a linear behavior $T^{2}\Gamma_{\bot}\propto 2+\phi_{\bot}+(1+3\phi_{\bot})\left(\frac{T}{T_{\rm N}}-1\right)$ and is a decreasing function of the temperature for $\phi_{\bot}<-1/3$. Using Eq. (31), it corresponds to $\delta\lambda/\lambda>1/2$. The threshold value $\delta\lambda/\lambda=1/2$ is plotted in Fig. 3 (b) as a dotted gray line. For values $\delta\lambda>\lambda/2$, the curve of $T^{2}\Gamma_{\bot}$ lies above this limiting case (dotted gray line), becomes non-monotonic, and exhibits a maximum.

### v2

@@ III Equations of motion
In this Section, starting from the FP equation shown in Eq. (16), we derive the equation of motion for the spin polarization, identified as the first moment [11], and for the two-time spin-spin correlation function [30, 26].
@@ III.1 Spin polarization
where on the right-hand side the first line represents the precession term, the second the diffusion term, and the third the Gilbert damping term. For the sake of clarity, we exclude the inter-sublattice Gilbert term $\Lambda^{\prime}$ from the present discussion. The full derivation of the equation of motion of the spin polarization, including $\Lambda^{\prime}$, is provided in Appendix A. Note that the anisotropy terms $\delta D_{\rm T}$ and $\delta\lambda$ only affect the in-plane component of the magnetization, and that the evaluation of the time derivative of the first-order moment involves higher-order moments (namely, second and third). The difficulty of solving the FP equation is transferred to the generation of an infinite hierarchical set of equations, which involves the higher-order moments. To solve this, in analogy to the approach used in [11] for a ferromagnetic classical model, we apply the MF approximation. The PDF $\mathcal{P}_{\mathrm{MF}}$ assumes that the spins of the sublattices, ${\bf\it s}_{A}$ and ${\bf\it s}_{B}$, are independent. Each one is coupled to an effective external field $k_{\rm B}T{\bf\it\xi}_{\mathrm{A}}$ and $k_{\rm B}T{\bf\it\xi}_{\mathrm{B}}$, respectively, where $k_{\rm B}T$ is the thermal energy  [11]. The PDF $\mathcal{P}_{\mathrm{MF}}$ is expressed as
@@
where $B(\xi)$ denotes the Langevin function [28]
@@
$\dot{\bm{m}}_{\alpha}$ $=\bm{m}_{\alpha}\times\expectationvalue{{\bm{h}}_{\alpha}}_{\rm MF}$ (22)
@@
Eq. (24) coincides with the nonlinear equation that determines the global minimum of the free energy via the classical Hamiltonian, as expressed in Eq. (5), within the MF approximation [18]. Proceeding with the assumption that there is a stationary solution with spins lying in the $x$-$y$ plane, e.g. along the $x$ direction such that ${\bf\it\xi}_{\mathrm{A}}=-{\bf\it\xi}_{\mathrm{B}}=\xi{\bf\it u}_{x}$, the resulting expression for $\xi$ is given by
@@ III.2 Spin-spin correlation functions
where the indices $i$ and $\ell$ refer to the Cartesian coordinates $x,y,z$, $\mathcal{P}({\bf\it\mathcal{S}},t_{1}|{\bf\it\mathcal{S}}^{\prime},t_{0})$ is the conditional probability of having the spin configuration ${\bf\it\mathcal{S}}=({\bf\it s}_{\mathrm{A}},{\bf\it s}_{\mathrm{B}})$ at time $t_{1}$ assuming that the configuration at time $t_{0}$ was ${\bf\it\mathcal{S}}^{\prime}=({\bf\it s}^{\prime}_{\mathrm{A}},{\bf\it s}^{\prime}_{\mathrm{B}})$, $P_{0}({\bf\it\mathcal{S}}^{\prime})$ is the stationary probability distribution [30]. Since we enforce the AFM condition $\bm{m}_{A}=-\bm{m}_{B}$ with $\Delta>0$, and we set ${\bf\it\xi}_{\mathrm{A}}=-{\bf\it\xi}_{\mathrm{B}}=\xi{\bf\it u}_{z}$. In the stationary regime, the correlation functions depend only on the time difference, so that $g_{\alpha i,\beta\ell}(t_{1},t_{0})\to g_{\alpha i,\beta\ell}(t)$, with $t=t_{1}-t_{0}$ for simplicity.
By differentiating Eq. (30) with respect to time through the use of Eq. (16), we derive the equation governing the dynamics of the correlation functions. The complete derivation is detailed in Appendix B, where it is shown that achieving an exact resolution of the motion equation requires solving a hierarchical series of equations. Since this hierarchy is infinite, an exact solution is unfeasible. Therefore, to derive a closed expression for the correlation functions, it becomes necessary to truncate the hierarchy. In the spirit of the Gaussian approximation, the truncation is performed by assuming that Wick’s theorem is applicable to the correlation functions of fluctuations of order higher than second [18], allowing them to be expressed in terms of second-order correlation functions. Initial conditions are fixed considering that in the limit $t\to 0$, we have $\mathcal{P}({\bf\it\mathcal{S}},t|{\bf\it\mathcal{S}}^{\prime},0)\rightarrow\delta({\bf\it\mathcal{S}}-{\bf\it\mathcal{S}}^{\prime})$. It follows that
@@ III.2.2 In-plane dynamics
which is expressed in terms of the zero-temperature spin-wave energy gap $\Omega_{0}=\sqrt{\Delta(2J+\Delta)}$  [44], and also incorporates a renormalization contribution stemming from the damping correction $\Gamma_{\bot}^{\prime}$. For the sake of clarity, the inter-sublattice terms $\lambda^{\prime}$ and $\delta\lambda^{\prime}$ have been omitted here. They have been examined in detail in Appendix B.2, where it is shown how they are incorporated into $\Gamma_{\bot}$ and $\Gamma_{\bot}^{\prime}$.
@@ IV Applications
Employing the method used to derive the equations of motion for spin polarization and spin-spin correlation functions, we examine two potential applications. Firstly, we investigate the dynamics of the spin-wave generated within the AFM phase [24]. Secondly, we show a phenomenological model tailored to describe the effects of magnetic fluctuations on the experimentally observed resistance fluctuations in MPX3 [15].
@@ IV.1 Spin-waves
Starting from the results of Sec. III.1, we derive the equation of motion for spin polarization in the AFM configuration being perturbed by a weak, external magnetic field $\bm{h}^{\rm ext}$ to both sublattices [13].
@@ IV.2 Thermal fluctuations in the resistance
As mentioned in the introduction, this work is motivated by the experimental observations presented in [15], where vertical resistance fluctuations in a 2D antiferromagnetic semiconductor showed an anomalous behavior close below the Néel temperature. The compound used in the experiment is a transition metal-phosphorus-trichalcogenides (MPX3[5, 36, 15, 45]. In these compounds, transition metal atoms (M) lay on a plane forming a honeycomb lattice, with negligible orbital overlap that inhibits conduction electrons from hopping within the plane, electron transport, however, takes place principally via the phosphorus-tricalchogenide groups (PX3) that take the shape of pyramids above and below the metal layer, and mediate longitudinal charge transport as sketched in Fig. 2 [5, 36, 15, 45].
@@
In this framework, one can approximate the electrons responsible for conduction as distinct from those responsible for the magnetic properties [51, 5, 36]. Based on the definition of the internal magnetic field for the sublattice in Eq. (7), and in the macrospin approximation, we can define the net internal magnetic field as the sum of the two fields ${\bf\it h}_{A}$ and ${\bf\it h}_{B}$:
@@
In the presence of the exact AFM configuration $\bm{s}_{A}=-\bm{s}_{B}$, the conducting electrons are not affected by a net internal magnetic field. When a disturbance disrupts the perfect AFM order, a non-zero net internal field can be produced. Using the methods shown in Sec. III, we develop a phenomenological model to describe thermal fluctuations as the source of this disturbance. According to the scheme represented in Fig. 2, it is assumed that transport, due to conducting electrons in a (stray) magnetic field, generates an electrical resistance that depends quadratically on the intensity of the internal net magnetic field $\bm{h}$ as
$R=R_{0}+\frac{\kappa_{R}}{2}h^{2}+\frac{\kappa_{R}^{\prime}}{2}h_{z}^{2}~,$ (54)
where $R_{0}$, $\kappa_{R}$, and $\kappa_{R}^{\prime}$ denote phenomenological parameters. Here, $R_{0}$ is the resistance in the absence of any interaction with the AFM system, $\kappa_{R}$ is a proportionality constant that is taken to be independent of the magnetic configuration, and $\kappa_{R}^{\prime}$ is a similar parameter that can additionally account for a cylindrical symmetry in this effective magnetoresistance. The underlying physical picture is that the electrons responsible for charge transport are distinct from those determining the magnetic behavior, similarly to the model for antiferromagnetic metals described in Ref. [51]. We proceed in analogy with conducting systems subjected to a weak external magnetic field, for which the electrical resistance usually exhibits a quadratic dependence on the intensity of the applied magnetic field [21], including the cases with vertical tunneling transport [33, 22, 10]. Accordingly, within the present framework, the resulting internal net magnetic field can affect the electrical resistance. We therefore treat the influence of the effective net field as a perturbation and assume that the resistance can be systematically expanded in powers of this field. In addition, the 2D system is assumed to be invariant under spatial inversion, as in the minimal description of MPX3 compounds on a honeycomb lattice. As a result, the lowest-order nonzero terms in the perturbative series are quadratic in the effective net field. Finally, because we use a 2D honeycomb lattice as our reference model (with analogous results applying to a 2D square lattice), the system is rotationally invariant about the $z$-axis in the long-wavelength limit [8]. This symmetry reduces the description to only two phenomenological coefficients, $\kappa_{R}$ and $\kappa_{R}^{\prime}$. For simplicity, we take the resistance to depend spherically on the net effective magnetic field, which is equivalent to setting $\kappa_{R}^{\prime}=0$.
With the application of the FP equation, the resistance $R(t)$ is considered a function of the random variables $\bm{s}_{\alpha}(t)$, linked to the internal net magnetic field $\bm{h}(t)$. It is useful to define the effective magnetoresistance as $R(t)=R_{0}+\kappa_{R}[R_{\parallel}(t)+R_{\bot}(t)]$, where
@@
$\expectationvalue{R}=R_{0}-\kappa_{R}\Big[\Delta(2J+\Delta)\frac{B(\xi)}{\xi}+(J+\Delta)^{2}B^{2}(\xi)\Big]~,$ (56)
where the terms in square brackets are the contributions in- and out-of-plane, respectively. Since we are interested in the resistance correlation function $\expectationvalue{\delta R(t)\,\delta R(0)}$, we write $\delta R(t)=R(t)-\expectationvalue{R}=\kappa_{R}[\delta R_{\parallel}(t)+\delta R_{\bot}(t)]$, where
@@
Our goal is to calculate the resistance power spectrum [49, 25, 7], defined as
@@
Fig. 3 (a) shows the out-of-plane component $S_{\parallel}(\omega)$ as a function of the temperature, covering the range from zero to Néel temperature, and setting $\Delta/J=0.35$. In particular, we consider three values of $\omega$: $\omega=\lambda J/10$ (solid red line), $\omega=\lambda J$ (solid green line) and $\omega=20\lambda J$ (solid blue line). Moreover, for each $\omega$, as reference, the dashed line and the dotted line represent, respectively, the contributions in Eq. (61) that involve the damping rate $\Gamma_{\parallel}^{\prime}$, which is $\propto 4\sigma_{z}^{2}B^{2}(\xi){\cal L}(\omega,\Gamma^{\prime}_{\parallel})+\sigma_{z}^{4}{\cal L}(\omega,2\Gamma_{\parallel}^{\prime})$ and $\Gamma_{\parallel}$, which is $\propto\sigma_{z}^{4}{\cal L}(\omega,2\Gamma_{\parallel})$. Here, we observe that in $S_{\parallel}(\omega)$, the term involving $\Gamma_{\parallel}$ (dotted lines) exhibits a monotonic behavior with temperature arising from the increase of thermal fluctuations. Specifically, as the temperature increases, there is an enhancement of the prefactor $\sigma_{z}^{4}$. On the other hand, the terms of $S_{\parallel}(\omega)$ involving $\Gamma_{\parallel}^{\prime}$ exhibit a peak at a temperature lower than $T_{\rm N}$, and the specific value depends on the chosen $\omega$. In particular, unlike $\sigma_{z}$, the staggered magnetization decreases with increasing temperature, thus the interplay between $\sigma_{z}$ and $B(\xi)$ results in the non-monotonic behavior and the formation of a peak. Moreover, it has been observed that when $\omega\lesssim\lambda J$, the terms involving $\Gamma_{\parallel}^{\prime}$ play a dominant role. As a result, the entire $S_{\parallel}(\omega)$ displays a distinct peak at a temperature lower than $T_{\rm N}$. Fig. 3 (b) shows the behaviour of the characteristic damping rates $\Gamma_{\parallel}$ (black solid line) and $\Gamma_{\parallel}^{\prime}$ (black dashed line) as functions of temperature. The horizontal lines represent the values of $\omega$: $\lambda J/10$ and $\lambda J$, used in Fig. 3 (a), serving as references. Within the explored temperature range, $\Gamma_{\parallel}^{\prime}$ vanishes at Néel temperature, while $\Gamma_{\parallel}$ remains finite. This panel provides a quantitative criterion for determining when the term associated with $\Gamma_{\parallel}$ becomes negligible, that is, for $\omega<\Gamma_{\bot}$.
@@
Figure 4: a) In-plane term $S_{\bot}(\omega)$ of the resistance power spectrum, defined in Eq. (65), shown as a function of the temperature between the zero temperature and the Néel temperature, for different values of $\omega$: $\omega=\lambda J/10$ (red lines), $\omega=\lambda J$ (green lines), and $\omega=20\lambda J$ (blue lines). Here, the dimensionless parameter $\delta\lambda$ is set $\delta\lambda=\lambda$ (solid lines), $\delta\lambda=0$ (dashed lines), and $\delta\lambda=-\lambda/2$ (dash-dotted lines). We obtain a non monotonic behavior only for $\omega=20\lambda J$ and $\delta\lambda=\lambda$. b) $T^{2}\Gamma_{\bot}$ as a function of temperature, where the damping rate is defined in Eq. (37a). The association between line style and $\delta\lambda$ remains consistent with that in panel a), while the gray dotted line represents $\delta\lambda=\lambda/2$. Results in all panels have been obtained by setting $\Delta=0.35J$.
@@
which is proportional to the Lorentzian function centered at zero and of width $2\Gamma_{\bot}$ multiplied by $\sigma_{\bot}^{4}$. Since we are interested in temperatures below $T_{\rm N}$, we replace $\sigma_{\bot}^{4}=(k_{\rm B}T)^{2}/(J+\Delta)^{2}$. Using Eq. (37a), we see that considering $\omega$ in units of $\lambda J$ makes $S_{\bot}(\omega)$ in the form of a product of $1/\lambda$ and a function of $\omega/(\lambda J)$ and $\delta\lambda/\lambda$. Again, this allows the results to be shown without explicit dependence on $\lambda$.
@@
Recalling that $\delta\lambda\neq 0$ implies that the diffusion coefficients in Eq. (15b) and the Gilbert damping parameters are different, we note that for $\omega\lesssim\lambda J$ when $\delta\lambda>0$, the diffusion term and damping in planar spin-dynamics become more efficient compared to the cases with $1\leq\delta\lambda/\lambda<0$. The lower bound of $\delta\lambda/\lambda=-1$ ensures that diffusive coefficients and the Gilbert damping terms remain non-negative. Except for the case with $\omega=20\lambda J$ and $\delta\lambda=-1$, which will be discussed below, the in-plane contribution to the resistance power spectrum exhibits a monotonic temperature dependence, as expected. Here, increasing temperature improves fluctuations, which in turn leads to an increase in the prefactor $\sigma_{\bot}^{4}\propto T^{2}$.
In Fig. 4 (a), this case is shown by the solid blue line that displays a peak at a temperature slightly below the Néel temperature (black triangle). The reason for this behaviour lies in the relation between $\omega$ and $\Gamma_{\bot}$. In particular, if $\omega\gg\Gamma_{\bot}$ across the entire temperature range examined, it is reasonable to approximate ${\cal L}(\omega,2\Gamma_{\bot})\approx\Gamma_{\bot}/(\pi\omega^{2})$, as a consequence of Eq. (65) one has $S_{\bot}(\omega)\propto T^{2}\Gamma_{\bot}/\omega^{2}$. The quantity in the numerator is shown in Fig. 4 b), where the different line styles are encoded in the same way as in Fig. 4 a), where we see that for $\delta\lambda/\lambda=1$ (solid line) $T^{2}\Gamma_{\bot}$ exhibits a peak below $T_{\rm N}$. Using Eqs. (28) and (37a), one finds that as $T\to 0$, the quantity $T^{2}\Gamma_{\bot}$ exhibits an increasing quadratic temperature dependence, since $\Gamma_{\bot}\approx(\lambda+\delta\lambda)(J+\Delta)$. On the other hand, as $T\to T_{\rm N}^{-}$, by using $\xi\propto\sqrt{1-T/T_{\rm N}}$, it shows a linear behavior $T^{2}\Gamma_{\bot}\propto 2+\phi_{\bot}+(1+3\phi_{\bot})\left(\frac{T}{T_{\rm N}}-1\right)$ and is a decreasing function of the temperature for $\phi_{\bot}<-1/3$. Using Eq. (38), it corresponds to $\delta\lambda/\lambda>1/2$. The threshold value $\delta\lambda/\lambda=1/2$ is plotted in Fig. 4 (b) as a dotted gray line. For values $\delta\lambda>\lambda/2$, the curve of $T^{2}\Gamma_{\bot}$ lies above this limiting case (dotted gray line), becomes non-monotonic, and exhibits a maximum.